A book containing 100 pages numbers from 1 to 100. a range is opened at random and is selected. Find the probability the opened page is multiple of 6 or 10


1
Expert's answer
2021-08-16T16:50:23-0400

Total number of pages: "N=100"

Favorable pages:


"6,10,12,18,20,24,30,36,40,42,48,50,"

"54,60,66,70,72,78,80,84, 90, 96, 100"

Total number of favorable pages: "n=23"


"\dfrac{100}{6}=16+\dfrac{4}{3}"

"30, 60, 90"

"n=16+10-3=23"

"p=\dfrac{n}{N}=\dfrac{23}{100}=0.23"